A consumer organisation wants to obtain information about , the mean number of drawing pins in the boxes of a certain brand which, according to the label, should contain 100 pins. In nine randomly chosen boxes this organisation finds the following numbers of drawing pins:

90 94 88 92 90 86 94 90 86



(a) (i) Test whether < 100. Take as level of significance a= 0:05.

(ii) Which assumption do you need?

(b) If the true value of u is 101, what type of error did you make in part (a)?

(c) (i) Determine a 90% confidence interval for u

(ii) Is it likely that u= 101?
1
Expert's answer
2021-07-13T03:10:33-0400

Part (a)

(i)

"t=\frac{r\sqrt{n-2}}{\sqrt{1-r^2}}\\
t=\frac{(-0.9594)\sqrt{9-2}}{\sqrt{1-0.9594^2}}\\
t=-8.9996\\

df=n-2=9-2=7\\
(1-\alpha)=0.95\\
\alpha=0.05\\
\frac{\alpha}{2}=0.025\\"

The corresponding probability of 0.975 is -2.306

Test statistic is "-8.9996< t_s-2.306"

Ho is rejected


(ii) The population form which sample is drawn is normally distributed


(b) A Type II error (i.e. do not reject the null hypothesis H0 when it was in fact false)

(c) (i)

"90+-Z_{1-0.05}*\frac{6}{\sqrt{9}}\\
90+-1.645*\frac{6}{3}\\
90+-1.96*2\\
(86.71,93.29)"

(ii)Yes